OpenStudy (anonymous):

how to transform this dy/dt-y=t+1 to it"s general sol?

OpenStudy (anonymous):

this is linear eq

OpenStudy (anonymous):

One thing I would so is just guess that it is a polynomial. \[ y = At+B \\ y' = A \\ y'-y = A - (At+B) = -At + A - B \\ y'-y = t+1 = -At + A-B \implies (1)t + (1) = (-A)t + (A-B) \]Then try to solve for our coefficients: \[ 1 = -A \\ 1 = A-B \]

OpenStudy (anonymous):

Next, we want to find the solution in the homogeneous case: \[ y' - y = 0 \]For first order differential equations, we will be in the form: \[ y' + f(t) y = 0 \]And the homogeneous solution is just: \[ e^{\int f(t)dt} \]

OpenStudy (anonymous):

So add up the homogeneous and non-homogeneous parts, to get: \[ Ce^{t}-t-2 \]

OpenStudy (anonymous):

I meant to write \[ e^{-\int f(t)dt} \]

OpenStudy (anonymous):

@wio why -ve there at the integrating facto?should be positive right?

OpenStudy (anonymous):

Oh, yeah, it should be positive, I was mixing up: \[ y' +f(t) y= 0 \]with \[ y' = f(t)y \]

OpenStudy (anonymous):

\[ \large \mu (t) = e^{\int-dt} = e^{-t} \]\[ \large \begin{array}{rc'} e^{-t}y' - e^{-t}y &=& e^{-t}(t+1) \\ (e^{-t}y)' &=& te^{-t} + e^{-t} \\ e^{-1}y &=& (-te^{-t}-e^{-t}) + (-e^{-t}) \\ y &=& -t - 2 \end{array} \]

OpenStudy (anonymous):

Weird............

OpenStudy (foolaroundmath):

dy/dt = y+t+1 y+t+1 = v dv/dt = dy/dt+ 1 => dy/dt = dv/dt -1 replacing, we get dv/dt -1 = v dv/dt = v+1 dv/(v+1) = dt integrate to get ln(v+1) = t +c ln(y+t+1) = t+c y+t+1 = ke^t y = ke^t -t -1

OpenStudy (anonymous):

@FoolAroundMath thanks for the help, but one thing doesn't add up: \[ y = ke^t-t-1 \\ y' = ke^t-1 \\ y' - y = (ke^t-1) - (ke^t-t-1) = ke^t-1 - ke^t + t + 1 = t = t \neq t+1 \]

OpenStudy (foolaroundmath):

yeah my bad... made a mistake upon substitution: reposting dy/dt = y+t+1 y+t+1 = v dv/dt = dy/dt+ 1 => dy/dt = dv/dt -1 replacing, we get dv/dt -1 = v dv/dt = v+1 dv/(v+1) = dt integrate to get ln(v+1) = t +c ln(y+t+2) = t+c y+t+2 = ke^t y = ke^t -t -2